A population consists of four numbesr (3, 8, 10, 15). Consider all possible sample of size 2 that can be drawn without replacement from the population.

Find the following:

a. Population Mean

b. Population Variance

c. Population Standard Deviation

d. The Mean of the sampling distribution of means

e. The Standard deviation of the sampling distribution of means


1
Expert's answer
2021-10-29T02:52:16-0400

a.

 We have population values "3,8,10,15," population size "N=4"

"\mu=\dfrac{3+8+10+15}{4}=9"

b.


"\sigma^2=\dfrac{1}{4}((3-9)^2+(8-9)^2+(10-9)^2"

"+(15-9)^2)=18.5"


c.


"\sigma=\sqrt{\sigma^2}=\sqrt{18.5}\approx4.3012"


d. We have population values "3,8,10,15," population size "N=4" and sample size "n=2." Thus, the number of possible samples which can be drawn without replacement is



"\dbinom{N}{n}=\dbinom{4}{2}=6""\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 3,8 & 5.5 \\
 \hdashline
 2 & 3,10 & 6.5 \\
 \hdashline
 3 & 3,15 & 9 \\
 \hdashline
 4 & 8,10 & 9 \\
 \hdashline
 5 & 8,15 & 11.5 \\
 \hdashline
 6 & 10,15 & 12.5 \\
 \hline
\end{array}"

The sampling distribution of the sample means.



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
& \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 & 5.5 & 1 & 1/6 & 11/12 & 121/24 \\
 \hdashline
 & 6.5 & 1& 1/6 & 13/12 & 169/24 \\
 \hdashline
 & 9 & 2 & 1/3 & 3 & 27 \\
 \hdashline
 & 11.5 & 1& 1/6 & 23/12 & 529/24 \\
 \hdashline
 & 12.5 & 1 & 1/6 & 25/12 & 625/24 \\
 \hdashline
 Total & & 6 & 1 & 9 & 523/6 \\ \hline
\end{array}"




"E(\bar{X})=\sum\bar{X}f(\bar{X})=9.8"




"\mu=\dfrac{6+8+10+12+13}{5}=9.8"

The mean of the sampling distribution of the sample means is equal to the the mean of the population.



"E(\bar{X})=\mu_{\bar{X}}=9.8=\mu"



e.


"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"




"=\dfrac{523}{6}-(9)^2=\dfrac{37}{6}"




"\sigma_{\bar{X}}=\sqrt{Var(\bar{X})}=\sqrt{\dfrac{37}{6}}\approx2.4833"

Verification:


"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{18.5}{2}(\dfrac{4-2}{4-1})"




"=\dfrac{37}{6}, True"